OpenStudy (anonymous):

Solve 2x + y = 7 x – y = –4

OpenStudy (anonymous):

add bouth equations: 2x + y = 7 x – y = –4 ----------- + 3x =3 x=1 1-y=-4 y=5

OpenStudy (tennistar):

given. 2x + y = 7 ...............1 x - y = - 4 =>x = y - 4 ................2 Now put this value of x in equation 1. 2 ( -4 + y ) + y =7 =>-8 + 2y + y = 7 =>3y =15 =>y = 5 Now put this value of y in equation 2. x= 5 - 4 =>x = 1 (x,y) = (1,5)

OpenStudy (anonymous):

then how woul i do this one ?? What is a possible first step to eliminate a variable in the following system? 7x – 9y = 2 3x – 6y = –4

OpenStudy (tennistar):

<<7x-9y=2 <<3x-6y=-4 When doing elimination the idea is to get one of the variables to cancel out. 2(7x-9y)=2(2) = 14x-18y=4 [first equation] -3(3x-6y)=-3(-4) = -9x+18y=-12 [second equation] You'll notice that the y's in this case cancel each other out. Multiply the first equation by 2 and the second equation by -3 <<<<Answer

OpenStudy (anonymous):

thats what i got too(: